# Twisted Fourier transforms on non-Kac compact quantum groups

### v1

@@ 5.2. The case of non-Kac and non-coamenable free orthogonal quantum groups
To demonstrate $[0,1]\subsetneq I(O_{F}^{+},p)$, our strategy is to apply a recently developed Haagerup inequality on non-Kac free orthogonal quantum groups [BVY21, You22] to establish the following lemma.
@@
Let $\mathbb{G}$ be a compact quantum group satisfying $\left\|Q_{\alpha}\right\|_{op}=\left\|Q_{\alpha}^{-1}\right\|_{op}$ for all $\alpha\in\text{Irr}(\mathbb{G})$. If $x\in I(\mathbb{G},p)$ with $21\leq p<$ 2, then we have $\sup_{\alpha\in\text{Irr}(\mathbb{G})}\frac{\left\|Q_{\alpha}\right\|_{op}^{|2x-1|}}{d_{\alpha}}<\infty$.
@@
$[(Q_{\alpha})_{jj}^{-1}(Q_{\alpha})_{ii}]^{(\frac{1}{p}-\frac{1}{2})x}\cdot\left(\frac{(Q_{\alpha})_{ii}^{-1}}{d_{\alpha}}\right)^{\frac{1}{p}}\leq K\cdot\left\|f\right\|_{L^{2}(\mathbb{G})}=K\cdot\sqrt{\frac{(Q_{\alpha})_{ii}^{-1}}{d_{\alpha}}}.$ (5.33)
@@
For the first inclusion, it is clear that $[0,1]\subseteq I(O_{F}^{+},p)$ by Theorem 3.6, so let us suppose that $x\notin[0,1]$, i.e. $|2x-1|>1$. Then, by Lemma 5.4, we obtain
@@
Thus, we can conclude that $x\in I(O_{F}^{+},1)$, and this implies $x\in I(O_{F}^{+},p)$ for all $21\leq p<$ 2 by Theorem 3.6.
(2) Conversely, for the second inclusion, let us suppose $x\in I(O_{F}^{+},p)$ with $21\leq p<$ 2. Then, Lemma 5.5 implies
@@
The following Corollary is a direct consequence of Theorem 5.6, demonstrating that $[0,1]\subsetneq I(\mathbb{G},p)$ can occur within the category of non-Kac and non-coamenable compact quantum groups.
@@
Let $O_{F}^{+}$ be a non-Kac free orthogonal quantum group. If $\left\|F\right\|_{op}^{6}<\frac{d_{1}+\sqrt{d_{1}^{2}-4}}{2}$, then $[0,1]\subsetneq I(O_{F}^{+},p)$ holds for all $21\leq p<$ 2.
@@
$21<|2x-1|<\frac{\log\left(\frac{d_{1}+\sqrt{d_{1}^{2}-4}}{2}\right)}{2\log\left\|F\right\|_{op}}-$ 2 . (5.48)
@@
The condition $\left\|F\right\|_{op}^{6}<\frac{d_{1}+\sqrt{d_{1}^{2}-4}}{2}$ in Corollary 5.7 can be true only for non-coamenable free orthogonal quantum groups. Indeed, for non-Kac coamenable free orthogonal quantum groups, we may assume that $O_{F}^{+}$ is the twisted quantum group $SU_{q}(2)$ with $10<q<$ 1, and $F$ is canonically chosen to be $\left[\begin{array}[]{cc}0&-q^{\frac{1}{2}}\\ q^{-\frac{1}{2}}&0\end{array}\right]$. Then, the condition $\left\|F\right\|_{op}^{6}<\frac{d_{1}+\sqrt{d_{1}^{2}-4}}{2}$ is equivalent to
@@
Acknowledgements: The author was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (No.RS-2025-00561391 and No.RS-2024-00413957) and by Samsung Science and Technology Foundation under Project Number SSTF-BA2002-01.

### v2

@@ 5.2. The case of non-Kac and non-coamenable free orthogonal quantum groups
To demonstrate $0,1]\subsetneq I(O_{F}^{+},p)$, our strategy is to apply a recently developed Haagerup inequality on non-Kac free orthogonal quantum groups [BVY21, You22] to establish the following lemma.
@@
Let $\mathbb{G}$ be a compact quantum group satisfying $\left\|Q_{\alpha}\right\|_{op}=\left\|Q_{\alpha}^{-1}\right\|_{op}$ for all $\alpha\in\text{Irr}(\mathbb{G})$. If $x\in I(\mathbb{G},p)$ with $1\leq p<2$, then we have $\sup_{\alpha\in\text{Irr}(\mathbb{G})}\frac{\left\|Q_{\alpha}\right\|_{op}^{|2x-1|}}{d_{\alpha}}<\infty$.
@@
$(Q_{\alpha})_{jj}^{-1}(Q_{\alpha})_{ii}]^{(\frac{1}{p}-\frac{1}{2})x}\cdot\left(\frac{(Q_{\alpha})_{ii}^{-1}}{d_{\alpha}}\right)^{\frac{1}{p}}\leq K\cdot\left\|f\right\|_{L^{2}(\mathbb{G})}=K\cdot\sqrt{\frac{(Q_{\alpha})_{ii}^{-1}}{d_{\alpha}}}.$ (5.33)
@@
For the first inclusion, it is clear that $0,1]\subseteq I(O_{F}^{+},p)$ by Theorem 3.6, so let us suppose that $x\notin[0,1]$, i.e. |2x1|>1|2x-1|>1. Then, by Lemma 5.4, we obtain
@@
Thus, we can conclude that $x\in I(O_{F}^{+},1)$, and this implies $x\in I(O_{F}^{+},p)$ for all $1\leq p<2$ by Theorem 3.6.
(2) Conversely, for the second inclusion, let us suppose $x\in I(O_{F}^{+},p)$ with $1\leq p<2$. Then, Lemma 5.5 implies
@@
The following Corollary is a direct consequence of Theorem 5.6, demonstrating that $0,1]\subsetneq I(\mathbb{G},p)$ can occur within the category of non-Kac and non-coamenable compact quantum groups.
@@
Let $O_{F}^{+}$ be a non-Kac free orthogonal quantum group. If $\left\|F\right\|_{op}^{6}<\frac{d_{1}+\sqrt{d_{1}^{2}-4}}{2}$, then $0,1]\subsetneq I(O_{F}^{+},p)$ holds for all $1\leq p<2$.
@@
$1<|2x-1|<\frac{\log\left(\frac{d_{1}+\sqrt{d_{1}^{2}-4}}{2}\right)}{2\log\left\|F\right\|_{op}}-2.$ (5.48)
@@
The condition $\left\|F\right\|_{op}^{6}<\frac{d_{1}+\sqrt{d_{1}^{2}-4}}{2}$ in Corollary 5.7 can be true only for non-coamenable free orthogonal quantum groups. Indeed, for non-Kac coamenable free orthogonal quantum groups, we may assume that $O_{F}^{+}$ is the twisted quantum group $SU_{q}(2)$ with $0<q<1$, and $F$ is canonically chosen to be $\left[\begin{array}[]{cc}0&-q^{\frac{1}{2}}\\ q^{-\frac{1}{2}}&0\end{array}\right]$. Then, the condition $\left\|F\right\|_{op}^{6}<\frac{d_{1}+\sqrt{d_{1}^{2}-4}}{2}$ is equivalent to
@@
6. Application to a similarity problem for $L^{1}(\mathbb{G})$
A Banach algebraic formulation of the well-known similarity problem for locally compact groups asks whether for every bounded algebra homomorphism $\Psi:L^{1}(G)\rightarrow B(H)$ there exists an invertible operator $T\in B(H)$ such that $T\circ\Psi(\cdot)\circ T^{-1}$ is a *-homomorphism, where $L^{1}(G)$ is the convolution algebra of $G$. This problem admits a natural extension to the setting of locally compact quantum groups [BDS13]. In this broader framework, however, the objects of study are completely bounded algebra homomorphisms of $L^{1}(\mathbb{G})$, rather than bounded ones, reflecting the perspective of operator space theory. Here, the convolution product $\star$ on $L^{1}(\mathbb{G})$ is given by
$(\varphi\star\psi)(a)=(\varphi\otimes\psi)(\Delta(a)),~a\in L^{\infty}(\mathbb{G}).$ (6.1)
Motivated by this perspective, the construction of explicit examples of bounded algebra homomorphisms of $L^{1}(\mathbb{G})$ that fail to be completely bounded has emerged as an important problem. To the best of our knowledge, the only known examples arise from the duals of non-abelian free groups [CS13] and reduced free products of compact quantum groups [BDS13]. Aside from the reduced free product technique, no examples of bounded but non-completely bounded algebra homomorphisms of $L^{1}(\mathbb{G})$ are currently available.
[…]
$\left[\pi^{(0)}(\varphi)\right](\alpha)=(\text{id}\otimes\varphi)\left(u^{\alpha}\right)=\left(\varphi(u^{\alpha}_{ij})\right)_{1\leq i,j\leq n_{\alpha}}$ (6.2)
for all $\alpha\in\text{Irr}(\mathbb{G})$. A strong connection between $\pi^{(0)}$ and our standard Fourier transform $\mathcal{F}^{(0)}_{1}$ is described by
$\mathcal{F}^{(0)}_{1}(\varphi)=\pi^{(0)}(\varphi\circ S),~\varphi\in L^{1}(\mathbb{G}),$ (6.3)
where $S$ is the antipode map. In particular, if $\mathbb{G}$ is a compact group $G$, then (6.3) reduces to
$\mathcal{F}^{(0)}_{1}(f)=\pi^{(0)}(\check{f}),~f\in L^{1}(G),$ (6.4)
where $\check{f}(x)=f(x^{-1})$ for all $x\in G$.
While $\pi^{(0)}$ is slightly different from our standard Fourier transform $\mathcal{F}^{(0)}_{1}$, an advantage of this approach is an application to the study of the similarity property of the convolution algebra $L^{1}(\mathbb{G})$. For any $x\in\mathbb{R}$ and $\varphi\in L^{1}(\mathbb{G})$, let us consider
$\pi^{(x)}(\varphi)=Q^{\frac{x}{2}}\pi^{(0)}(\varphi)Q^{-\frac{x}{2}}\in\prod_{\alpha\in\text{Irr}(\mathbb{G})}M_{n_{\alpha}}(\mathbb{C})$ (6.5)
where $Q=(Q_{\alpha})_{\alpha\in\text{Irr}(\mathbb{G})}$.
Proposition 6.1.
[…]
Proof.
First, let us focus on the extremal case $x=1$. Then, as in the proof of Theorem 3.5, note that $(Q_{\alpha}^{\frac{1}{2}}\otimes 1)(u^{\alpha})^{c}(Q_{\alpha}^{-\frac{1}{2}}\otimes 1)$ is a unitary for any $\alpha\in\text{Irr}(\mathbb{G})$. Thus,
$\left[(Q_{\alpha}^{\frac{1}{2}}\otimes 1)(u^{\alpha})^{c}(Q_{\alpha}^{-\frac{1}{2}}\otimes 1)\right]^{*}=\sum_{i,j=1}^{n_{\alpha}}Q_{\alpha}^{-\frac{1}{2}}e^{\alpha}_{ji}Q_{\alpha}^{\frac{1}{2}}\otimes u^{\alpha}_{ij}$ (6.6)
is also a unitary element in $M_{n_{\alpha}}(\mathbb{C})\otimes L^{\infty}(\mathbb{G})$. This implies
$\left\|(\text{id}\otimes\varphi)\left(\sum_{i,j=1}^{n_{\alpha}}Q_{\alpha}^{-\frac{1}{2}}e^{\alpha}_{ji}Q_{\alpha}^{\frac{1}{2}}\otimes u^{\alpha}_{ij}\right)\right\|_{M_{n_{\alpha}}}\leq\left\|\varphi\right\|_{L^{1}(\mathbb{G})}.$ (6.7)
Since the left-hand side is given by
$\sum_{i,j=1}^{n_{\alpha}}Q_{\alpha}^{-\frac{1}{2}}e^{\alpha}_{ji}Q_{\alpha}^{\frac{1}{2}}\cdot\varphi(u^{\alpha}_{ij})=Q_{\alpha}^{-\frac{1}{2}}[\pi^{(0)}(\varphi)](\alpha)^{t}Q_{\alpha}^{\frac{1}{2}}.$ (6.8)
and the transpose map is contractive on $M_{n_{\alpha}}(\mathbb{C})$, we obtain
$\left\|\pi^{(1)}(\varphi)\right\|_{\ell^{\infty}(\widehat{\mathbb{G}})}=\sup_{\alpha\in\text{Irr}(\mathbb{G})}\left\|Q_{\alpha}^{\frac{1}{2}}[\pi^{(0)}(\varphi)](\alpha)Q_{\alpha}^{-\frac{1}{2}}\right\|_{M_{n_{\alpha}}}\leq\left\|\varphi\right\|_{L^{1}(\mathbb{G})}.$ (6.9)
Then, the contractivity of $\pi^{(x)}$ for arbitrary $0\leq x\leq 1$ follows from the complex interpolation method as in the second step of the proof of Theorem 3.5.
Second, for any $\varphi,\psi\in L^{1}(\mathbb{G})$ and $\alpha\in\text{Irr}(\mathbb{G})$,
$\left[\pi^{(0)}(\varphi\star\psi)\right](\alpha)=(\text{id}\otimes(\varphi\star\psi))(u^{\alpha}).$ (6.10)
If we write $u^{\alpha}=\sum_{i,j=1}^{n_{\alpha}}e^{\alpha}_{ij}\otimes u^{\alpha}_{ij}$, then the right-hand side is given by
$(\text{id}\otimes\varphi\otimes\psi)(\text{id}\otimes\Delta)(u^{\alpha})=\sum_{i,j=1}^{n_{\alpha}}\sum_{k=1}^{n_{\alpha}}e^{\alpha}_{ij}\cdot\varphi(u^{\alpha}_{ik})\cdot\psi(u^{\alpha}_{kj})$ (6.11)
$=\sum_{i,j=1}^{n_{\alpha}}\sum_{k=1}^{n_{\alpha}}e^{\alpha}_{ij}\cdot\left[\pi^{(0)}(\varphi)\right](\alpha)_{ik}\left[\pi^{(0)}(\psi)\right](\alpha)_{kj}$ (6.12)
$=\left[\pi^{(0)}(\varphi)\right](\alpha)\left[\pi^{(0)}(\psi)\right](\alpha)$ (6.13)
[…]
$\pi^{(x)}(\varphi\star\psi)=Q^{\frac{x}{2}}\pi^{(0)}(\varphi\star\psi)Q^{-\frac{x}{2}}=Q^{\frac{x}{2}}\pi^{(0)}(\varphi)\pi^{(0)}(\psi)Q^{-\frac{x}{2}}$ (6.14)
$=Q^{\frac{x}{2}}\pi^{(0)}(\varphi)Q^{-\frac{x}{2}}\cdot Q^{\frac{x}{2}}\pi^{(0)}(\psi)Q^{-\frac{x}{2}}=\pi^{(x)}(\varphi)\pi^{(x)}(\psi),$ (6.15)
[…]
Let us focus on the special case $x=1$. Then the contractive algebra homomorphism $\pi^{(1)}$ is not completely bounded in general by the following theorem.
Theorem 6.2.
Let $\mathbb{G}$ be a non-Kac compact quantum group satisfying
$\sup_{\alpha\in\text{Irr}(\mathbb{G})}\frac{n_{\alpha}\max\left\{\left\|Q_{\alpha}\right\|,\left\|Q_{\alpha}^{-1}\right\|\right\}}{d_{\alpha}}=\infty$ (6.16)
Then the contractive algebra homomorphism $\pi^{(1)}$ is not completely bounded.
Proof.
By [Pis03, Theorem 2.5.2], if we assume that $\pi^{(1)}$ is completely bounded, then there exists $V\in\ell^{\infty}(\widehat{\mathbb{G}})\overline{\otimes}L^{\infty}(\mathbb{G})$ such that
$(\text{id}\otimes\varphi)(V)=\pi^{(1)}(\varphi)=Q^{\frac{1}{2}}\pi^{(0)}(\varphi)Q^{-\frac{1}{2}}$ (6.17)
for all $\varphi\in L^{1}(\mathbb{G})$. This implies that, if we write $V=(V(\alpha))_{\alpha\in\text{Irr}(\mathbb{G})}$ with $V(\alpha)\in M_{n_{\alpha}}(\mathbb{C})\otimes L^{\infty}(\mathbb{G})$, then we should have
$V(\alpha)=(Q_{\alpha}^{\frac{1}{2}}\otimes 1)u^{\alpha}(Q_{\alpha}^{-\frac{1}{2}}\otimes 1).$ (6.18)
Then, it is straightforward to see that
$(\text{id}\otimes h)(V(\alpha)V(\alpha)^{*})$ $=\frac{n_{\alpha}}{d_{\alpha}}Q_{\alpha}^{-1},$ (6.19)
$(\text{id}\otimes h)(V(\alpha)^{*}V(\alpha))$ $=\frac{n_{\alpha}}{d_{\alpha}}Q_{\alpha},$ (6.20)
and this implies $\left\|V(\alpha)\right\|^{2}\geq\frac{n_{\alpha}\max\left\{\left\|Q_{\alpha}\right\|,\left\|Q_{\alpha}^{-1}\right\|\right\}}{d_{\alpha}}$. Hence, the given assumption (6.16) implies the following contradiction
$\left\|V\right\|_{\ell^{\infty}(\widehat{\mathbb{G}})\otimes L^{\infty}(\mathbb{G})}^{2}=\sup_{\alpha\in\text{Irr}(\mathbb{G})}\left\|V(\alpha)\right\|_{M_{n_{\alpha}}\otimes L^{\infty}(\mathbb{G})}^{2}=\infty.$ (6.21)
Example 6.3.
In particular, $\pi^{(1)}$ is not completely bounded for all non-Kac free orthogonal quantum groups $O_{F}^{+}$ since the sufficient condition (6.16) is satisfied. Here, $F\in GL_{N}(\mathbb{C})$ such that $F\overline{F}=\pm\text{Id}_{N}$, $\text{Irr}(O_{F}^{+})\cong\mathbb{N}_{0}=\left\{0,1,2,\cdots\right\}$ and $\left\|Q_{k}\right\|=\left\|Q_{k}^{-1}\right\|$ holds for all $k\in\mathbb{N}_{0}$. If $N=2$, then $L=\inf_{k\in\mathbb{N}_{0}}\frac{\left\|Q_{k}\right\|}{d_{k}}>0$, so it is straightforward to see
$\sup_{k\in\mathbb{N}_{0}}\frac{n_{k}\left\|Q_{k}\right\|}{d_{k}}\geq\left(\sup_{k\in\mathbb{N}_{0}}n_{k}\right)\cdot L=\infty.$ (6.22)
If $N\geq 3$, an exponential growth of $\frac{n_{k}\left\|Q_{k}\right\|}{d_{k}}$ is proved in the proof of [BVY21, Theorem 3.3], so this implies
$\sup_{k\in\mathbb{N}_{0}}\frac{n_{k}\left\|Q_{k}\right\|}{d_{k}}=\infty.$ (6.23)